WHAT VOLUME OF 0.10M h2so4 must be add to 50 ml of 0.6 M NAOHsolution to make a solution in which the molarity of the h2so4 is 0.050m?

Grade:11

2 Answers

ishan jogalekar
6 years ago
0.1V + 50×0.6=0.05×(0.6+V)0.1V+3=0.03+0.05Vso,V=59.4mlhere we have taken 50ml of naoh of 0.6molar in V volume of h2so4 of0.10molar to make h2so4 of o.o5molar sol..so m1v1+m2v2=m3v3.so ans. is 59.4ml
Rohit
6 years ago
We can use M1V1=M2V2 i.e., 10/100*V1=6/10 *50/1000 . On solving , V1=0.30L *1000(Converted in ml). Final answer is 300ml